# Attention as an Adaptive Filter

### v1

@@ 2.6 Robust Maximum Likelihood Trajectory Estimation
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. If our dynamic model, defined by $\{\mathbf{S},\mathbf{S}^{-1},\bm{\Lambda},\bm{\Lambda}_{C},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}\}$, does not perfectly match reality, these theoretical precision matrices may misrepresent the true uncertainty, leading to suboptimal state estimates.
One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big{(}0,\mathbf{V}^{C}_{ij}\big{)}$. We generally can’t observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big{(}0,\mathbf{V}^{C}_{ij}+\mathbf{R})$. Note that these pair-wise residuals are distinct from the total innovation $\mathbf{y}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\bar{z}}_{ij}$. Since $\mathrm{Cov}\!\left(\mathbf{y}_{i}\right)=\left(\sum_{j}\mathbf{P}^{C}_{ij}\right)^{-1}\preceq\big{(}\mathbf{P}^{C}_{ij}\big{)}^{-1},\,\forall j$, the combined estimator’s innovation covariance is no larger than the residual covariance of any single predictor.
2.6.1 Variational Formulation of Robust Estimation
Suppose we have vector-valued measurements:
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\geq\mathcal{L}(q,\bm{\mu}_{i}):=\mathbb{E}_{q(\mathbf{\Sigma}_{ij})}\big{[}\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big{]}-\mathrm{KL}\big{(}q(\mathbf{\Sigma}_{ij})\,\|\,p(\mathbf{\Sigma}_{ij})\big{)}.$
$\mathbf{\Sigma}_{ij}^{-1}=\tau_{ij}\,\mathbf{P}^{C}_{ij},$
$\tau_{ij}\sim\mathrm{Gamma}(\alpha,\beta).$
The expected precision under the posterior is:
where:
The corresponding (marginal) likelihood $p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})$ is a Student’s $t$. This is equivalent to an M-estimator with loss
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
Using the fact that $\log(1+x)\approx x\quad\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
2.6.2 Iteratively-Reweighted Least Squares
As we noted above, we cannot observe the true latent state $\bm{\mu}_{i}$, so we cannot compute the residual $\mathbf{r}_{ij}:=\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}$. However, we can use our current best estimate of $\bm{\mu}_{i}$, which we denote $\mathbf{\hat{z}}_{ii}$, and a corresponding residual: $\mathbf{\hat{r}}_{ij}:=\mathbf{\hat{z}}_{ij}-\mathbf{\hat{z}}_{ii}$. This leads naturally to an iteratively-reweighted least squares (IRLS)-style update for the latent state (at iteration $k$):
$\mathbf{\hat{z}}^{(k+1)}_{ii}=\Big{(}\sum_{j}w_{ij}^{(k)}\mathbf{P}^{C(k)}_{ij}\Big{)}^{-1}\sum_{j}w_{ij}^{(k)}\mathbf{P}^{C(k)}_{ij}\mathbf{\hat{z}}_{ij}^{(k)},$
Recall that (by construction) $\mathbf{P}_{ij}^{C(k)}=\mathbf{S}^{-\dagger}\bm{\lambda}_{\mathbf{P}ij}^{C}\mathbf{S}^{-1}$. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}=e^{-\bm{\Lambda}\Delta t_{ij}}$, define a corresponding residual $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$, and let $\bm{\lambda}_{\mathbf{P}ij}^{C}$ be the diagonal of $\bm{\lambda}_{\mathbf{P}ij}^{C}$. We have an iterative update rule for the state estimate in the eigenbasis:
@@ 2.6.2 Iteratively-Reweighted Least Squares
$w_{ij}^{(k)}=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{C})^{\top}\|\mathbf{r}^{(k)}_{\mathbf{s}ij}\|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. (Alternatively, if we only have a single iteration, we can replace $\bm{\lambda}_{\mathbf{P}ij}^{C}$ with $\bm{\lambda}_{\mathbf{P}ij}^{C}+\bm{\lambda}_{R}$ (the covariance of $\mathbf{z}_{\mathbf{s}i}$).) To stabilize the iteration, we can instead use an affine combination of the current and previous estimate:
@@
This IRLS iteration, with precision matrices derived from an LTI model, performs a batched maximum likelihood update of latent states using pairwise uncertainty propagation. We will refer to it as Kalman-IRLS, as it is closely related to adaptive Kalman Filtering, although unlike the latter, it does not rely on recursive updates. It can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\hat{z}}^{(k)}_{ii}\rightarrow\mathbf{z}^{C}$ in expectation with increasing $k$, we can take $\mathbf{\bar{z}}_{\mathbf{s}}\approx\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}$ at the last layer, and then predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{\text{pred}}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}$
Finally, we compute the output prediction: $\mathbf{z}^{\text{pred}}_{i}=\mathbf{S}\mathbf{z}^{\text{pred}}_{\mathbf{s}i}$, which we then compare to the actual measurement at the next time step $\mathbf{z}_{i+1}$. We can then define a loss, e.g. $\mathcal{L}=\sum_{i}\|\mathbf{z}^{\text{pred}}_{i}-\mathbf{z}_{i+1}\|_{2}^{2}$ and train our model parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ using backpropagation.
2.6.3 Iteratively-Reweighted Kalman Filter (IRKF)
Prediction Step (Include Process Noise).
$\mathbf{\bar{z}}_{\mathbf{s}i}^{-}=\bm{\Lambda}\mathbf{\bar{z}}_{\mathbf{s},i-1}^{+},\quad\bm{\Lambda}_{\mathbf{V},i}^{-}=\bm{\Lambda}\bm{\Lambda}_{\mathbf{V},i-1}^{+}\bm{\Lambda}^{\dagger}+\bm{\Lambda}_{Q}=\|\bm{\Lambda}\|^{2}\bm{\Lambda}_{\mathbf{V},i-1}^{+}+\bm{\Lambda}_{Q},\quad\bm{\Lambda}_{\mathbf{P},i}^{-}=(\bm{\Lambda}_{\mathbf{V},i}^{-})^{-1}.$
Weighted Update (Measurement + IRKF Iteration).
Within each iteration $k$ at time $t_{i}$, compute the innovation:
Define the nominal posterior precision from the measurement model:
and compute a residual-based robust weight:
The weighted posterior precision is then:
Notice that the posterior information vector can be written as:
So the total posterior mean in the eigenbasis becomes:
Alternatively, we can use a damped update:
@@ 3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the batch MLE form of the above adaptive filter using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{k}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{k}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{k}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
@@
We can define separate key/query and value dynamics parameters: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{k}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. If we set $d_{k}=d_{v}$, then we can use a shared set of parameters for the key and value dynamics: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\bm{\lambda}_{C}^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
@@
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}\in\mathbb{C}^{d_{k}\times m},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}\in\mathbb{C}^{d_{k}\times m},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}\in\mathbb{C}^{d_{v}\times m}$
Exponential tensors: $\mathcal{E}_{\mathbf{qk}}\in\mathbb{C}^{d_{k}\times m\times m},\quad\mathcal{E}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{E}_{\mathbf{qk}}[k,i,j]=e^{\lambda_{\mathbf{qk}k}(t_{i}-t_{j})},\quad\mathcal{E}_{\mathbf{v}}[k,i,j]=e^{\lambda_{\mathbf{v}k}(t_{i}-t_{j})}$
"Pulled-forward" estimates: $\mathcal{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{d_{k}\times m\times m},\quad\mathcal{\hat{Z}}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
@@
The residuals: $\mathcal{R}\in\mathbb{C}^{d_{k}\times m\times m}$:
@@
$\|\mathcal{R}_{\mathbf{qk}}[k,i,j]\|^{2}=\mathrm{Re}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)+\mathrm{Im}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)$
The key and value precision tensors: $\mathcal{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{d_{k}\times m\times m},\quad\mathcal{P}^{C}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
@@
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\|\mathcal{R}_{\mathbf{qk}}[k,i,j]\|^{2}/\nu\right)^{-1}$
@@
After $n$ layers with shared weights, compute the prediction $\mathbf{Z}^{\text{pred}}\in\mathbb{C}^{d_{e}\times m}$ using a one-step forward transition:
$\mathbf{Z}_{\mathbf{v}}^{\text{pred}}[k,i]=\mathcal{E}_{\mathbf{v}}^{\text{pred}}[k,i+1,i]\cdot\mathbf{\bar{Z}}[k,i]$
@@
$\mathbf{Z}^{\text{pred}}[k,i]=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{\text{pred}}[k,i]$
This adaptive filter generalizes attention by introducing a more structured similarity measure that integrates uncertainty and time decay. If $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{\mathbf{k}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}}\rightarrow\mathbf{0}$, the relative temporal decay on the attention weights disappears and we are left with a tensor generalization of standard attention.
By using the same (diagonal) matrix exponential to filter and to predict the next step, we enforce self-consistency in the model. One issue is that it may be preferable to do multiple steps of iterative refinement of the estimated trajectory before prediction. One way to address this is to give the model extra flexibility by outputting a gated interpolation between estimate and prediction. That is:
@@
$\mathbf{g}^{(k)}=\sigma\!\left(\mathbf{a}^{\top}\phi\!\left(\|\mathbf{r}^{(k)}\|^{2}\right)+\mathbf{b}\right),$

### v2

@@ 2.3 Robust Maximum Likelihood Trajectory Estimation
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual has distribution $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}\big)$. We generally can’t observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}+\mathbf{R})$. We can then rescale the precision matrices with an adaptive weight:
$\mathbf{P}^{C}_{ij}\leftarrow\mathbf{P}^{C}_{ij}/w_{ij},\quad\text{where: }\,w_{ij}=\frac{1}{1+\mathbf{r_{z}}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r_{z}}_{ij}/\nu},$
where $\nu\in\mathbb{R}$ is a scaling factor. This leads naturally to an iteratively-reweighted least squares (IRLS) iteration in the eigenbasis. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}=e^{-\bm{\Lambda}\Delta t_{ij}}$, and define a corresponding measurement residual in the eigenbasis $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$. Let $\bm{\lambda}_{\mathbf{P}ij}^{C}=\text{diag}(\bm{\Lambda}_{\mathbf{P}ij}^{C})$. We apply the iterative update:
@@
$w_{ij}^{(k)}=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{C})^{\top}\big|\mathbf{r}^{(k)}_{\mathbf{s}ij}\big|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. (Alternatively, if we only have a single iteration, we can replace $\bm{\lambda}_{\mathbf{P}ij}^{C}$ with $\bm{\lambda}_{\mathbf{P}ij}^{C}+\bm{\lambda}_{R}$ (the covariance of $\mathbf{z}_{\mathbf{s}i}$).) To stabilize the iteration, we can instead use an affine combination of the current and previous estimate (i.e. a residual connection):
@@
This IRLS iteration, with precision matrices derived from an LTI model, can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\hat{z}}^{(k)}_{ii}\rightarrow\mathbf{z}^{C}$ in expectation with increasing $k$, we can take $\mathbf{\bar{z}}_{\mathbf{s}}\approx\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}$ at the last layer, and then predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{+}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}$
Finally, we compute the output prediction: $\mathbf{z}^{+}_{i}=\mathbf{S}\mathbf{z}^{+}_{\mathbf{s}i}$, which we then compare to the actual measurement at the next time step $\mathbf{z}_{i+1}$. We can then define a loss, e.g. $\mathcal{L}=\sum_{i}\big|\mathbf{z}^{+}_{i}-\mathbf{z}_{i+1}\big|_{2}^{2}$ and train our model parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ using backpropagation.
@@ 3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the batch MLE form of the above adaptive filter using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{qk}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
@@
We can define separate key/query and value dynamics parameters: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{qk}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. If we set $d_{qk}=d_{v}$, then we can use a shared set of parameters for the key and value dynamics: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\big|\bm{\lambda}_{C}\big|^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
@@
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}\in\mathbb{C}^{d_{qk}\times m},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}\in\mathbb{C}^{d_{qk}\times m},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}\in\mathbb{C}^{d_{v}\times m}$
Exponential tensors: $\mathcal{E}_{\mathbf{qk}}\in\mathbb{C}^{d_{qk}\times m\times m},\quad\mathcal{E}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{E}_{\mathbf{qk}}[k,i,j]=e^{\bm{\lambda}_{\mathbf{qk},k}(t_{i}-t_{j})},\quad\mathcal{E}_{\mathbf{v}}[k,i,j]=e^{\bm{\lambda}_{\mathbf{v},k}(t_{i}-t_{j})}$
"Pulled-forward" estimates: $\mathcal{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{d_{qk}\times m\times m},\quad\mathcal{\hat{Z}}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
@@
The residuals: $\mathcal{R}\in\mathbb{C}^{d_{qk}\times m\times m}$:
@@
$\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}=\mathrm{Re}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)+\mathrm{Im}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)$
The key and value precision tensors: $\mathcal{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{d_{qk}\times m\times m},\quad\mathcal{P}^{C}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
@@
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}/\nu\right)^{-1}$
@@
After $n$ layers with shared weights, compute the prediction $\mathbf{Z}^{+}\in\mathbb{C}^{d_{e}\times m}$ using a one-step forward transition:
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=\mathcal{E}_{\mathbf{v}}[k,i+1,i]\cdot\mathbf{\bar{Z}}[k,i]$
@@
$\mathbf{Z}^{+}[k,i]=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{+}[k,i]$
This adaptive filter generalizes attention by introducing a more structured similarity measure that integrates uncertainty and time decay. By using the same (diagonal) matrix exponential to filter and to predict the next step, we enforce self-consistency in the model. One issue is that it may be preferable to do multiple steps of iterative refinement of the estimated trajectory before prediction. One way to address this is to give the model extra flexibility by outputting a gated interpolation between estimate and prediction. That is:
@@
$\mathbf{g}^{(k)}=\sigma\!\left(\mathbf{a}^{\top}\phi\!\left(\big|\mathbf{r}^{(k)}\big|^{2}\right)+\mathbf{G}\right),$